Accelerometer suspended in a vehicle braking from 50km/HR to 0 in a distance of 60 m. What angle does it make with vertical?
1
Expert's answer
2020-10-29T10:16:28-0400

As per the question,

Initial speed "(u)=50km/hr = 50\times\frac{1000}{3600}m/sec=\frac{125}{9}m/s"

Final speed "(v)=0"

Covered distance "(d)=60m"

Let the acceleration is "=a"

Now, "v^2=u^2-2ad"

Now, substituting the values,

"\Rightarrow a=\frac{u^2}{2d}=\frac{125^2}{2\times 9^2\times 60}m/s^2"

"\Rightarrow a=1.60 m/s^2"